OpenStudy (anonymous):

Given that the following points lie on y= f(x), find the coordinates of the point each moves to under the transformation y=3f(2x). i) (3,-5) ii) (1,2) iii (-2,1)

OpenStudy (rob1525):

y=f(x) and x=2x so... y=3*-5=-15 and x= 2*3 =6. New point = (6,-15). Do the same for the others.